### v1

# Why Human Guidance Matters in Collaborative Vibe Coding

Abstract
Writing code has been one of the most transformative ways for human societies to translate abstract ideas into tangible technologies. Modern AI is transforming this process by enabling experts and non-experts alike to generate code without actually writing code, but instead, through natural language instructions, or “vibe coding”. While increasingly popular, the cumulative impact of vibe coding on productivity and collaboration, as well as the role of humans in this process, remains unclear. Here, we introduce a controlled experimental framework for studying collaborative vibe coding and use it to compare human-led, AI-led, and hybrid groups. Across 16 experiments involving 604 human participants, we show that people provide uniquely effective high-level instructions for vibe coding across iterations, whereas AI-provided instructions often result in performance collapse. We further demonstrate that hybrid systems perform best when humans retain directional control (providing the instructions), while evaluation is delegated to AI.
Keywords: Human-AI Collaboration; Complementarity; Vibe Coding; Large Language Models;
Figure 1: Vibe coding experimental paradigm. (A) Core procedure. An instructor views a reference image and its best SVG rendition from the previous iteration, and uses natural language instructions to guide code-generation. The Code Generator produces SVG code that can then be rendered into an image. (B) Iterated procedure. At each iteration, a selector chooses whether the current or the previous SVG image better matches the reference image. The selected SVG is then passed to an instructor, who provides vibe-coding instructions that are carried forward to the next iteration. (C) Interface of the human validation experiments. Participants rate the similarity of generated SVGs to the reference image.
Introduction
Coding is a foundational skill in contemporary digital society, supporting a wide range of economic, scientific, and creative activities [18, 14]. In recent years, advances in large language models have led to rapid growth in AI-assisted programming tools, accompanied by major investment from companies [43, 17]. Systems such as Copilot, Cursor, and related tools are now widely used in real-world development workflows [42, 19]. These tools do more than speed up coding and may even act as collaborative “thought partners” [7]. They may change how people approach problems, how tasks are structured, and how decisions are made when people build.
One new practice in human-AI collaborative coding is “vibe coding”. Vibe coding describes a conversational and intuition-driven style of programming wherein users provide high-level guidance and rely on iterative interaction with AI systems, rather than precise low-level control [31, 33, 10, 11, 8]. Instead of writing code, users describe goals, intentions, or desired changes, while AI systems generate and modify code accordingly. This interaction style reframes coding as a collaborative process between humans and AI, rather than a purely human activity.
These developments point toward the emergence of “hybrid societies” in which humans and AI agents jointly produce complex artifacts [34, 13]. In such systems, performance depends not only on the capabilities of individual AI models, but also on how humans and AI agents coordinate over time, how responsibilities are divided, and how guidance, evaluation, and execution are distributed across agents [41, 23, 33]. Despite the importance of these questions, we still have a limited empirical understanding of how AI-assisted coding reshapes collective behavior. Most prior work focuses on improving model accuracy, interface design, or single interaction usability, rather than the iterative process of collaboration and coordination in coding [22, 1, 2].
This gap persists in part because human–AI collaboration is difficult to study in the wild at scale [36]. Many studies rely on simulations of fully autonomous AI agents or offline benchmarks [40, 24], leaving open the question of whether direct human involvement – and what type of human involvement – changes outcomes in systematic ways. This raises two central questions: (i) is there a unique advantage to human–AI collaboration, rather than relying on fully automated AI pipelines?, and (ii) how should labor be effectively divided within groups that include both humans and AI?
To address these questions, we introduce a controlled experimental framework for studying collaborative vibe coding. Our approach affords full causal control over who provides guidance, who executes code, who validates outputs, and how information flows across iterations. We design a task that is intuitive for non-experts, supports repeated interaction, and allows for objective validation against a known target (see Figure 1). In this paradigm, we systematically compare human- and AI-driven vibe coding, analyze sources of misalignment, test hybrid collaboration strategies, and examine how different role allocations affect collective performance.
Our approach
Writing code allows people to translate arbitrary ideas into actions performed by machines. In this paper, we focus on one type of idea transfer, imagining a person wants to create a digital image. Here, provide a reference image as a stand-in for the idea people may want to communicate (Fig. 1A). The goal of the coding task is to materialize this visual idea through code. For simplicity, we use scalable vector graphics (SVG) as the coding medium. SVG provides a symbolic code that can be directly rendered into a visible image, allowing to easily judge whether the output matches the intended vision. This makes SVG well-suited for studying collaborative “vibe coding”, as participants can interpret the code by immediately observing its visual output. In this setting, a code instructor uses natural language to guide a code-generation system to produce SVG code that visually approximates the reference image. The process is iterative, emulating real-world coding practices: each instructor starts from the best result of the previous iterations and incrementally refines it to better match the target image.
Our goal is to determine under which conditions human–AI collaboration is most effective, and to assess what contribution people provide over fully automated AI coding systems.
Figure 2: Case study of human-led (human selectors and instructors) and AI-led (AI selectors and instructors) vibe coding. (A) Example progressions of the experiment for one reference image with human-led (top) and AI-led (bottom) vibe coding. (B) Examples from the last iteration of human-led (top) and AI-led (bottom) chains, more examples in Appendix Fig. 11 to 15.
Methods
Vibe coding experimental paradigm
Fig. 1A illustrate our setup (see screenshots from the experiment in Appendix Fig. 8B). An instructor (vibe coder) views a reference image and its SVG rendition from the previous iteration and then uses natural language instructions to guide a code-generation system. The code-generator produces SVG code that can then be rendered as an image. We created 10 reference images by prompting GPT-5 [35] to generate images of the following animals: cat, dog, tiger, bird, elephant, penguin, shark, zebra, giraffe, panda.
We also introduce a selection step to mimic realistic collaborative coding workflows. At each stage, the selector compares the current version with the previous one and can revert the changes if they do not improve the result. This ability to select and discard modifications reflects common processes of technical refinement, such as those used in version control systems (e.g., Git). Fig. 1B show the overall iterative procedure (see screenshots in Appendix Fig. 8C). At each iteration, a selector decides whether the current or the previous SVG better matches the reference image. The selected SVG is then passed to an instructor, who provides vibe-coding instructions that are passed forward.
In the basic experiment, the same participant performed both the selection and instruction roles at iteration $n$. Furthermore, in the first iteration, no previous SVG is provided, and the instructor generates an image based solely on the reference image. Likewise, in the second iteration, no selector is involved, as only a single image has been created at that point.
Finally, we simulate the role of independent evaluators by involving independent group of humans. These evaluators view both the reference image and the generated SVG output and assess the quality of the vibe-coding process by rating the visual similarity between the two images (Fig. 1C). Overall 120 people evaluated 4800 images (aggregated from all 16 experiments reported in this paper).
All experiments were implemented using PsyNet [12], an online framework for behavioral experiments. PsyNet naturally supports interactions between humans and AI agents within social networks [34, 26]. Appendix Fig. 16, and Appendix Fig 17 shows example of prompts provides to AI and human participants.
Participants and AI queries
Participants were recruited online via Prolific [30]. The participants were from the UK and identified English as their native language. We recruited 604 human participants overall. To simulate AI agents, we performed 4,800 API queries for code generation, and 5,327 API queries for instruction (Fig. 1A). The main AI model used was OpenAI GPT-5 [35], but for comparison we also performed additional 280 queries to Anthropic Claude Opus 4.5 [3] and 280 queries for Google Gemini 3 Pro [38]. Participants provided informed consent under the Cornell University–approved protocol IRB0148995, and were compensated at a rate of £10 per hour.
Results
Comparing human and AI roles
Figure 3: Performance of human- and AI-led vibe coding. Data points represent average rating scores across all experiments. Shaded area represents one standard error of the mean.
30 independent “chains” with 15 iterations were run (three repeats for the 10 reference images). We recruited 45 human participants, each of whom served as both selector and instructor across 10 chains (human-led vibe coding condition). Within each chain, the same participant first acted as the selector and then as the instructor. In a second condition (AI-led vibe coding), API calls to OpenAI’s GPT-5 model replaced human participants, while all other aspects of the experiment remained identical. In all cases, instructors and selectors used natural language to prompt a code-generation system (GPT-5).
Fig. 2A shows a representative example of multi-round vibe coding for one reference image. With humans as the selectors and instructors (Fig. 2A top), the SVG output improves steadily across iterations. In contrast, in the AI-led condition (Fig. 2A bottom), early iterations sometimes capture salient features, but later iterations do not substantially improve or even drift away from the target. Fig. 2B suggests a consistent human advantage in the final SVG outputs (15th iteration) across all target images under both conditions.
To test this quantitatively, we conducted validation experiments (Fig. 3; design described in Fig. 1C ). In the first iterations, people and AI systems attained similar ratings ($\Delta$=0.094, pp=1.000, Cohen’s dd=0.178), however in the final iteration humans significantly surpassed AI ($\Delta$=0.953, pp<0.001, Cohen’s dd=2.173). When human selectors and instructors were involved, similarity scores increased generally across iterations (indicated by a positive correlation rr=0.237, $\text{CI}_{95\%}$=[0.077, 0.387]) with highest relative improvement reaching 27.1%.
In contrast, in the AI-led baseline, similarity scores significantly decreased across iterations (indicated by a negative correlation rr=-0.351, $\text{CI}_{95\%}$=[-0.493, -0.183]). These results demonstrate a clear advantage of humans acting as selectors and instructors over a fully AI-driven coding pipeline. Moreover, while humans exhibit iterative refinement, AI systems deteriorate.
As we are using a single model in the AI-led condition while involving different participants in the human-led condition, the performance difference may be attributed to the collective intelligence of human groups. To test this, we recruited another 10 participants, each independently completing 3 chains (Asocial condition; Fig. 2C) . The performance was nearly identical, and increased continuously (rr=0.289, $\text{CI}_{95\%}$=[0.130, 0.436]). These results rule out individual differences in vibe coding ability as the primary driver of performance gains and suggest that improvement arises from alignment with the task objective rather than from social interaction per se. In other words, performance gains reflect human-guided iterative refinement and are largely independent of whether improvements are produced by a single individual or by a group. To examine whether the performance difference between AI and human instructors reflects a misalignment in their object representations [37], we conducted an additional experiment in which the evaluator role was replaced by multiple GPT-5 API calls (Fig. 10A). Specifically, we asked whether AI agents would recognize that their own outputs are inferior to those produced by humans, or instead show a preference for their own creations, which would indicate a potential alignment issue.
Figure 4: Comparing instruction semantics in human-led and AI-led conditions. (A) Example human- and AI-generated instructions. (B) UMAP projection of instructions in the embedding space and example word clouds of both human and AI instrubtions. (C) Validation experiment results from different AI-led experiments under instruction length limits. (D) Radar plot of seven semantic metrics: topic entropy (the diversity and unpredictability of topics), descriptive ratio (proportion of descriptive words), sentiment compound (overall emotion of the text), main IDF (how unique the vocabulary is), mean content length (average length of the answer), type token ratio (ratio of unique words to the total answer) and content ratio (proportion of content words vs. function words). For comparison, each metrics was normalized to the scale of 0 to 1.
We found a positive correlation between human and AI ratings (rr=0.348, pp<0.001; Fig. 10B) which is comparable to human and AI split halve reliability (humans: rr=0.340, AI: rr=0.448, both pp<0.001). However, AI evaluators assigned higher ratings to AI-generated outputs ($\beta$=0.298, SE=0.047, tt(895)=6.33, $\text{CI}_{95\%}$=[0.205, 0.390], pp<0.001). These findings suggest that the observed performance differences may stem from a misalignment in representations between humans and AI [15].
Figure 5: Hybrid Human-AI-Led vibe coding. (A) The schematics of human-AI hybrid vibe coding. (B) Validation rating results of human-AI hybrid vibe coding. (C) Trade-off between human proportion and the final vibe coding performance.
Divergences between human and AI instructions
To further explore what drove differences in the outcomes of human and AI instructions, we next analyzed the text of the instructions. Fig. 4A shows representative examples of human and AI instructions. Human instructions are typically short and goal-directed. In contrast, AI instructions are much longer and more complex. They often contain exhaustive descriptions of visual attributes, including fine-grained details about texture, lighting, color gradients, facial features, and animal-specific anatomical properties.
We next quantified semantic differences between human and AI instructions. We converted all instructions into sentence embeddings using GPT text-embedding-3-small and projected them into a two-dimensional space using UMAP ([27]; Fig. 4B). Human instructions form a single dense semantic cluster, indicating high consistency across targets and iterations. In contrast, AI instructions split into multiple target-specific clusters. This suggests that humans reuse a shared, task-oriented language across different targets, while AI treats each target animal image as a separate descriptive problem. Fig. 4B, show TF–IDF word clouds [21] demonstrating this phenomenon. Human word clouds are dominated by general action words such as “make”, “add”, “body”, and “head” which are reused across targets. AI word clouds contain many target-specific and highly descriptive terms. We then computed, across animal categories, the cosine similarity with average instruction embeddings. Human instructions showed higher similarities across categories (mean similarity M=0.841, $\text{CI}_{95\%}$=[0.828, 0.854]) compared with AI instructions (M=0.779, $\text{CI}_{95\%}$=[0.761, 0.794]). The difference was significant ($\Delta$=0.062, pp<0.001, Cohen’s dd=1.162).
AI instructions were much longer than human instructions (on average instruction length for humans was M=17.723, SD=16.166, and for AI was M=755.279, SD=193.803, pp<0.001 Cohen’s $d$ = 6.479). We next tested whether instruction length alone explains the performance gap. We reran the AI-led vibe coding condition with hard limits of 10, 20, or 30 words per instruction (Fig. 4C). After applying length limits, the mean content length of AI instructions became comparable to human instructions. However, limiting instruction length did not improve AI performance, continuing to show flat or declining performance over iterations (rr=-0.202, $\text{CI}_{95\%}$=[-0.372, -0.020], r=-0.219, $\text{CI}_{95\%}$=[-0.364, -0.068], r=-0.229, $\text{CI}_{95\%}$=[-0.382, -0.068] for 10, 20 and 30 iterations), indicating that verbosity alone does not account for the observed failure. Moreover, the semantic clustering of AI instructions remained substantially tighter than that of human instructions even when controlling for instruction length. For descriptions limited to 10, 20, and 30 words, AI instructions showed significantly smaller clustering than human instructions (within category similarity means M=0.639, 0.725, 0.756, respectively; compared with human’s M=0.396, all differences pp<0.001), indicating that AI remains overly target-specific even when constrained to short descriptions.
To further characterize the misalignment, we calculated seven semantic metrics for all instructions (Fig. 4D). After applying length limits, the mean content length of AI instructions became comparable to human instructions. However, AI instructions showed a higher descriptive ratio, indicating a greater proportion of descriptive words than action-oriented content. This suggests that AI focuses on detailed depiction of the target rather than on communicating effective guidance to the coding assistant. In addition, humans exhibited higher topic entropy than all AI conditions, indicating greater diversity and flexibility in how they approached improvements across iterations. This pattern suggests that humans are better at identifying what is missing or incorrect in the current SVG and at exploring alternative ways to fix it, whereas AI remains locked into narrow, over-specified descriptions.
Building on the previous results showing a strong advantage of humans as selectors and instructors, we next asked whether this advantage could be partially retained while reducing human involvement. Human labor is costly and difficult to scale, whereas AI can provide fast and generally comparatively less expensive automatic guidance. We next tested a hybrid human–AI vibe coding setting that mixes human and AI vibe coders within the same multi-round coding chain with the same and comparable code generation system.
Testing hybrid human-AI instructors and selectors
In this experiment, we constructed hybrid coding chains in which selector and instructor roles were randomly assigned to either a human or an AI agent according to a fixed ratio (Fig. 5A). As in all other experiments, human and AI instructors each interacted with an AI code generation system to create the SVG code. We tested three hybrid conditions: human-75% (AI-25%, where humans take 75% of iterations and AI takes the rest), human-50% (AI-50%), and human-25% (AI-75%). For each condition, we repeated the 10 target images twice, resulting in 20 coding chains and 300 total iterations. Since we found that the results were already stable with one iteration, each human or AI-bot participated in only one iteration. We recruited 225 human participants for the human-75% condition, 150 for the human-50% condition, and 75 for the human-25% condition.
As shown in Fig. 5B, all hybrid conditions performed better than fully AI-led vibe coding ($\Delta$=0.734, pp=0.007, Cohen’s dd=1.268, $\Delta$=0.350, pp=0.163, Cohen’s dd=0.736, $\Delta$=0.195, pp=0.571, Cohen’s dd=0.451 for human-75%, 50% and 25% conditions compared with the original AI condition), showing that even limited human involvement improves performance. However, performance decreased steadily as the proportion of AI increased: Human-75% condition (iterations 13-15: M=2.560, $\text{CI}_{95\%}$=[2.242, 2.994]; human-50% (M=2.176, $\text{CI}_{95\%}$=[1.921, 2.467], and human-25% conditions (M=2.021, $\text{CI}_{95\%}$=[1.780, 2.246]), showing a trade-off between the human proportion and the coding performance (rr=-0.986, pp=0.002; Fig. 5C). Moreover, none of the hybrid conditions showed significant improvement with iterations (correlations for human-75%, 50% and 25% were rr=0.043,-0.176, -0.23).
Role division in vibe coding
Beyond the question of whether vibe coding benefit from human-AI hybrid teams, it is an open question how roles should be allocated between different agents across iterations. In our framework, we distinguish between two roles – selection and instruction – reflecting common practices in real-world collaborative programming, where humans are involved both in choosing among alternatives and in providing guidance. To assess the respective contributions of these roles, we conducted an ablation study in which the selection step was removed and the edited output was always propagated to the next iteration.
We reran the experiment using 10 target images, repeated twice, yielding 20 coding chains and a total of 300 iterations. We recruited 30 new human participants to perform the task. As shown in Fig. 6A (light red), removing the selection step produced a significant reduction in human vibe-coding quality relative to the original human condition with selection in the final iteration (iterations 13-15: M=2.331, $\text{CI}_{95\%}$=[2.136, 2.552], which is significant compared with original human condition ($\Delta$=-0.448, pp=0.012, Cohen’s dd=-1.190). This shows that when humans guide the process, explicit evaluation and choice between alternatives helps maintain a coherent direction and supports cumulative improvement, though this effect is relatively small. However, for AI Fig. 6B (cyan), removing the selection had almost no effect (iterations 13-15: M=1.884, $\text{CI}_{95\%}$=[1.586, 2.233], compared with the original AI condition $\Delta$=0.058, pp=1.000, Cohen’s dd=0.112) suggesting that AI-generated guidance already follows a fixed internal preference, and explicitly comparing alternatives does not improve its trajectory.
Figure 6: Role division for collaborative vibe coding. (A) Validation results of experiments removing the code selection part. (B) Validation results of experiments setting human/AI agents in different roles (instructor/selector).
Fig 6B (light red) presents the results of a complementary ablation study with 69 participants in which the AI system assumed the selection role while humans provided the instruction. Replacing the human selector with an AI selector, while retaining a human instructor, yielded performance comparable to fully human-dominated vibe coding (iterations 13-15: M=2.748, $\text{CI}_{95\%}$=[2.552, 2.980], compared with the original human condition $\Delta$=-0.031, pp=1.000, Cohen’s dd=-0.080) suggesting that selection can be offloaded to AI with little loss in performance, offering a practical way to reduce human effort without sacrificing the quality of collaborative vibe coding.
This further suggests that human instructors tend to generate effective improvements. Moreover, human selection becomes essential when the instructor is an AI system (M=2.485, $\text{CI}_{95\%}$=[2.055, 2.930], compared with the original AI condition $\Delta$=0.659, pp=0.025, Cohen’s dd=1.055). Taken together, these findings highlight the importance of human supervision in mitigating compromised or suboptimal AI-generated outputs. Further analysis on selection acceptance rate (whether selector chooses the current SVG) proves this by revealing the misalignment between human and AI when collaborating on a task. Specifically, the acceptance rate when humans act as both instructor and selector (M=0.661, $\text{CI}_{95\%}=[0.620,0.700]$) is significantly higher than the condition with a human instructor and an AI selector (M=0.379, $\text{CI}_{95\%}=[0.334,0.423]$, difference $\Delta$=0.282, pp<0.001, Cohen’s dd=0.589) and vice versa (M=0.342, $\text{CI}_{95\%}=[0.295,0.388]$, $\Delta$=0.320, pp<0.001, Cohen’s dd=0.674).
Robustness to AI model type and social information
To test the robustness of our experimental vibe coding pipeline to other factors, we conducted several control experiments. First, we changed how the AI vibe coder evaluated the SVG outputs. In the main experiment, the AI vibe coder observed the SVG code directly. In the control conditions, we instead allowed the AI vibe coder to view the rendered SVG image only, matching the information available to human vibe coders, or to view both the SVG code and its rendered image. Across both viewing modes, AI-led vibe coding showed similar performance to the main AI condition, remaining significantly worse than human-led vibe coding ($\Delta$=-1.138, -0.545, both pp<0.01) and exhibiting no improvement across iterations (View SVG-rendered image: rr=-0.026, $\text{CI}_{95\%}$=[-0.174, 0.127]; View both SVG code and rendered image: rr=-0.321, $\text{CI}_{95\%}$=[-0.451, -0.181]; Fig. 7A). This indicates that the AI’s failure is not due to limited access to visual information.
We also tested the advanced method from previous research, named FeedBack Descent [20], where the selector not only provided the selection but also gave feedback on the code. Similarly, there is no improvement across iterations (rr=-0.07, $\text{CI}_{95\%}$=[-0.268, 0.110]) and the final performance is significantly lower than the human-led condition ($\Delta$=-0.677, pp=0.019, Cohen’s dd=-1.081; Fig. 7A).
We further tested whether the observed effects depend on the specific AI model used for vibe coding. In addition to GPT-5, we repeated the AI-led experiments using Claude-4.5-Opus (Anthropic, [3]) and Gemini-3-Pro (Google, [38]) as instructor and selector, while keeping the code generator to be GPT-5 (to facilitate a fair comparison between the AI model performance). Both alternative models produced lower final similarity scores than human-led vibe coding (Claude: M=2.323, $\text{CI}_{95\%}$=[1.979, 2.847]; Gemini: M=1.933, $\text{CI}_{95\%}$=[1.617, 2.324]) and showed the similar pattern of no improvement across rounds (Claude: rr=-0.046, $\text{CI}_{95\%}$=[-0.211, 0.117]; Gemini: rr=-0.129, $\text{CI}_{95\%}$=[-0.278, 0.026]; Fig. 7B). These results demonstrate that the performance gap and collapse effect are not specific to a single model, but instead reflect a general limitation of current AI systems when acting as high-level guidance providers in multi-round collaborative coding.
Figure 7: Control experiments for vibe coding pipeline. (A) Experiment results of different viewing modes in AI-led vibe coding. (B) Results of different AI-led vibe coding.
Discussion
Here, we systematically explore human-AI compositions and roles in vibe coding. Across multiple experiments, human-led coding consistently improved over iterations, while AI-led coding often collapsed despite access to the same information and similar execution capabilities. This points to key struggles of today’s AI systems in sustaining coherent high-level direction across repeated interactions, of the kind necessary for successful vibe coding. Our results highlight that studying human–AI collaboration directly reveals behaviors that would be missed by evaluating autonomous AI agents alone.
A key finding is the systematic misalignment between how humans and AI use language to guide collaboration. Humans tend to issue short, action-oriented instructions that focus on what to change next while AI produces long, descriptive instructions that emphasize details. Importantly, constraining instruction length did not eliminate this gap, indicating that misalignment arises from differences in strategy rather than verbosity. This can also be supported by our results that AI prefers its own creations rather than humans’. This aligns with broader concerns that language models may optimize for descriptive completeness instead of goal-directed coordination in interactive tasks [39, 28, 25].
Our hybrid and role-division experiments further clarify how labor should be allocated in human–AI systems. Partial human involvement improved performance relative to fully AI-driven pipelines, but even limited AI guidance could shift the coding trajectory in ways that later human input could not fully correct. At the same time, decomposing the vibe coder role showed that high-level idea generation and instruction are the critical human contributions, whereas evaluation and selection can often be delegated to AI without loss in performance. This suggests a practical design principle for hybrid systems: humans should set direction, while AI can support evaluation and execution.
More broadly, our findings contribute to a cognitive science of hybrid societies by showing that collective performance depends not only on agent competence, but on how roles, feedback, and control are assigned and structured over time [5]. Vibe coding offers a tractable model system for studying these dynamics, with implications beyond programming, toward other creative design work and decision-making in human–AI teams [2, 4, 32, 9, 7].
Limitations
Several limitations of our study should be noted. First, in our task, participants provided improvement instructions only after selecting the better SVG, whereas in real-world coding, people often also comment on the weaknesses of the rejected option; future work should incorporate comparative feedback on unselected outputs. Second, although SVG coding is intuitive and accessible, it remains a simplified task, and future studies should test more realistic coding settings, such as website or game development. Third, we used a single AI model as either a coding assistant, instructor, or selector, while prior work has explored mixtures of different AI systems; whether combining multiple AI agents improves vibe coding remains an open question. Finally, our experiments model collaboration as a linear iterative chain, whereas real-world coding often involves team-based, parallel collaboration, which should be explored in future work. We also note that we only study short-term human-AI interaction effects, but future work should also consider long-term feedback loops and risks like cognitive decline and overreliance [29, 16, 6].
Conclusion
To conclude, in this work, we introduced a controlled experimental framework for studying collaborative vibe coding and used it to systematically compare human- and AI-led vibe coding. We show that humans provide uniquely effective high-level guidance across iterations, that AI guidance often leads to performance collapse. Also, we find that a careful role allocation which keeps humans in charge of direction while offloading evaluation to AI can improve hybrid performance. Together, these results demonstrate the value of studying human–AI collaboration directly and offer concrete insights for designing effective hybrid human–AI systems.
Acknowledgments
This work was supported by the NSF grant “CollaborativeResearch: Research Infrastructure: HNDS-I: Building In-frastructure to Study Human-Al Hybrid Societies in Experi-mental Social Networks” (Award BCS-2523500) and partiallysupported by supported by the NSF grant “Collaborative Research:Designing smart environments to augment collectivelearning & creativity” (Award 171418)
ChatGPT version 5.2 (OpenAI) was used to assist manuscript editing and proofreading. Authors reviewed each of the edit suggestions, and approved the final version.
Appendix A Appendix
A. Experimental Interface
Here we present an example of the vibe coding experiment, which was also used as the teaching material in the experiment deployment.
Figure 8: Experimental interface. (A) Code Initialization. (B) Code editing as Fig. 1A. (C) Code selection and editing as Fig. 1B.
B. Validation Experiment Interface
Figure 9: Validation Rating Experiment Interface
C. Validation Results with AI as the evaluator
Figure 10: Performance and difference with AI as evaluator. (A) Validation performance when AI is the rating evaluator. (B) Relationship between rating scores of human evaluators and AI evaluators.
D. Qualitative Analysis of Vibe Coding
Here we provide more examples of vibe coding results extracted from the 15th iteration in the chain experiment. The first raw presents the reference images and the following presents different experimental conditions. In each subsection we all provide parts of examples from our two main human- and AI-led experiments as the baseline.
D.1 SVG Examples of Semantic Controlled AI-Led Experiments
Figure 11: Examples from all experiments - Part 1.
D.2 SVG Examples of Human-Led Asocial Experiment
Figure 12: Examples from all experiments - Part 2.
D.3 SVG Examples of Hybrid Human-AI Instructors & Selectors
Figure 13: Examples from all experiments - Part 3.
D.4 SVG Examples of Vibe Coding Role Division
Figure 14: Examples from all experiments - Part 4.
D.5 SVG Examples of AI Control Conditions
Figure 15: Examples from all experiments - Part 5.
E. System Prompts
Figure 16: System role of code generator
Figure 17: Experiment background for AI